OpenStudy (anonymous):

88% of what number is 66?

OpenStudy (paki):

88/100*x=66... solve for "x"....

OpenStudy (anonymous):

No

OpenStudy (anonymous):

66*100/88=?

OpenStudy (paki):

@OrangeMaster my explanation is rite or wrong....?

OpenStudy (anonymous):

66*100/88=75

OpenStudy (anonymous):

It's right @paki

OpenStudy (anonymous):

Why?

OpenStudy (paki):

hmm thanks @OrangeMaster

OpenStudy (anonymous):

you don't solve for x. cause I @cjackson2002 there is not an x @paki you added it in there which I don't understand why.

OpenStudy (anonymous):

can some one answer my question now please?

OpenStudy (paki):

just leave "x" alone at one side...and solve....

OpenStudy (paki):

my method is good explained....

OpenStudy (anonymous):

You shouldn't ask for help on someone else's post. Ask people in the chats.

OpenStudy (anonymous):

there was no x you had to just solve the question as it was written 66 times 100 divided by 88. okay sorry about me asking but how do I ask the class?